OpenStudy (anonymous):

i know how to prove it i think... but i dont understand what the professor means by this @ganeshie8 ||V+W||^2 = ||V||^2 + ||W||^2 + 2V*W (less than or equal to) ||V||^2 +||W||^2 +2||v||*||W|| = (||V|| + ||W||)^2 taking sqr ||V+W|| (less than or equal to) ||V||+||W||

OpenStudy (anonymous):

like i get how it equals the first and second "line" but then i dont get how he jumps to the final line like what happened to the +2||V||||W|| ?

ganeshie8 (ganeshie8):

\[V\bullet W = \|V\| \|W\| \cos\theta \]

ganeshie8 (ganeshie8):

would you agree \(V\bullet W \le \|V\| \|W\| \) ?

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good

OpenStudy (anonymous):

but how does that get into this "proof"?

ganeshie8 (ganeshie8):

so you agree that \(||V+W||^2 \le ||V||^2 +||W||^2 +2||v||*||W||\) ?

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

right hand side can be written as (||V|| + ||W||)^2 so thats same as \(||V+W||^2 \le (||V|| + ||W||)^2\) yes ?

OpenStudy (anonymous):

ok wait hold on... lol let me try to process it ._.

ganeshie8 (ganeshie8):

recall the formula \[\heartsuit^2+\spadesuit^2 + 2\heartsuit \spadesuit = (\heartsuit + \spadesuit)^2\]

OpenStudy (anonymous):

So i think i get. i got: ||V||^2 + ||W||^2 + 2V*W (less than or equal to) ||V||^2 + ||W||^2 + 2||V||||W|| so simply because 2V*W (less than or equal to) 2||V||||W|| it validates the upper equation, am i right?

ganeshie8 (ganeshie8):

Exactly!

OpenStudy (anonymous):

oh ok. so why after did he put... = (||V|| + ||W||) ^2 taking sqrt you get the answer like coming from MY equation how would taking the sqrt get to the answer? I just stopped there because i justified it but i never ended up back to ||V+W|| (less than or equal to) ||V||+||W||

ganeshie8 (ganeshie8):

\[||V+W||^2 \le (||V|| + ||W||)^2\] are you okays up to this step ?

OpenStudy (anonymous):

like i GET how it proves the inequality but how did he get rid of the 2V*W on the left side and 2||V||||W|| on the right to finally get ||V+W|| <or= ||V||+||W||

ganeshie8 (ganeshie8):

||V+W||^2 = ||V||^2 + ||W||^2 + 2V*W <= ||V||^2 + ||W||^2 + 2||V||||W || <= (||V|| + ||W||)^2

ganeshie8 (ganeshie8):

Notice that the left side is always ||V+W||^2 your prof is only messing witht he right hand side

OpenStudy (anonymous):

oh *palm to the face* ughhh thank you for being so patient with me :) i get it. ._.

ganeshie8 (ganeshie8):

haha np:)